please provide formulas with answer ​​​​​​

  1. The simple interest on a 45-day loan of $5,200 is $55.25. What is the simple interest APR?

  1. You borrowed $10,000 from your bank at 5.6% simple interest APR. You took out the loan on March 1 and plan on repaying it on August 1. What is the interest and maturity value of the loan?

  1. Franz is planning to borrow $182,000 from Columbia River Bank to build a road into his property. The road building (and loan) will commence on March 17th and the road will be completed on April 27th. At that time the property will be sold and the loan repaid to the bank (April 27th). Terms for the loan are simple interest at 2 (percentage) points over prime plus a 1.5% loan fee (the loan fee is added to the $ amount being borrowed). The Bank's prime rate is 8.75% APR. What is the maturity value?
  1. Bill is borrowing $150 from a pawnshop by giving the pawnshop a guitar worth $500. The pawnshop says she can repay the $150 plus 10% interest ($150 x 10 % = $15) in 30 days to redeem the guitar; if he does not repay the loan in 30 days, the pawn shop will sell his guitar. Help bill calculate the APR. (use 365 days)
  1. Lee is purchasing ski clothing from a supplier for $2,450. The seller offers a 4% discount if the invoice is paid within 10 days; if not paid within 10 days, the full amount must be paid within 30 days of the invoice date (4/10,n/30). Find the Annual Effective Rate. (use 365 days)
Answer #1

1.
=Interest/Loan*365/45
=55.25/5200*365/45
=8.6181%

2.
Maturity Value=Principal+Principal*rate*time
=10000+10000*5.6%*(153/365)
=10234.7397

Interest=10234.7397-10000=234.7397

3.
Maturity value=(loan+loan*loan fee%)*(1+(days/365)*rate)
=(182000+182000*1.5%)*(1+(41/365)*(8.75%+2%))
=186960.6780

4.
=Interest/Amount*365/days
=15/150*365/30
=121.6667%

5.
=(1/(1-discount%))^(365/days of no discount)-1
=(1/(1-4%))^(365/20)-1
=110.6444%